a watch cost 24.00 more than a calculator, and the calculator cosy 16.00 more than a game set.The total cost of the three items is 206.00.What is the cost of each item?Pls hel me answer this algebra problem, thanks.
Answer by htmentor(1343)

a watch cost 24.00 more than a calculator, and the calculator cosy 16.00 more than a game set.The total cost of the three items is 206.00.What is the cost of each item?
==================================
Let x = the cost of the game set
Then the cost of the calculator = 16 + x
And the cost of the watch = 24 + 16 + x = 40 + x
The formula for the total cost is:
x + 16 + x + 40 + x = 206
3x = 206 - 56
3x = 150
x = 50
So the cost of the game set is $50
The cost of the calculator is $66
The cost of the watch is $90